OpenStudy (anonymous):

Find the remainder using the remainder theorem: f(x) = 3x^4 - 3x^3 + x2 - 6x + 2 divided by (x + 7)

OpenStudy (raden):

x+7 = 0 x = -7 so, the remainder of f(x) when divided by (x+7) is f(-7) f(x) = 3x^4 - 3x^3 + x2 - 6x + 2 f(-7) = 3(-7)^4 - 3(-7)^3 + (-7)^2 - 6(-7) + 2 = ...

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

?

OpenStudy (raden):

can you show to me how you got -14 ??

OpenStudy (anonymous):

-2 times 7

OpenStudy (anonymous):

wait hold on

OpenStudy (anonymous):

probably mistake

OpenStudy (raden):

the start is here : 3(-7)^4 - 3(-7)^3 + (-7)^2 - 6(-7) + 2 = ...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

multiply?

OpenStudy (anonymous):

3(-7) = -21 ^4 =-84-3=-87

OpenStudy (raden):

to get good answer, do them one by one (-7)^4 = ... then times 3, get .... -3(-7)^3 = .... (-7)^2 = ... -6 * -7 = ... get the total aobve, then add 2

OpenStudy (anonymous):

so -84 for first

OpenStudy (anonymous):

second =63

OpenStudy (anonymous):

third -14

OpenStudy (anonymous):

last 42

OpenStudy (raden):

(-7)^4 = -7 x -7 x -7 x -7 = 2401

OpenStudy (anonymous):

then what

OpenStudy (raden):

3(-7)^4 = 3 * 2401 = 7203

OpenStudy (anonymous):

thats the answer 7203?

OpenStudy (raden):

it just for the first :p

OpenStudy (raden):

this's the 2nd -3 (-7)^3 = ... see that (-7)^3 = -7 x -7 x -7 = -343 so, -3(-7)^3 = -3 x -343 = 1029

OpenStudy (anonymous):

ok then?

OpenStudy (raden):

next (-7)^2 = -7 x -7 = 49

OpenStudy (raden):

next -6(-7) = 42

OpenStudy (anonymous):

yep

OpenStudy (raden):

see all calcution above is positive now find the total, f(-7) = 3(-7)^4 - 3(-7)^3 + (-7)^2 - 6(-7) + 2 = ... f(-7) = 7203 + 1029 + 49 + 42 + 2 = .......................

OpenStudy (anonymous):

-12,647

OpenStudy (anonymous):

the 1st

OpenStudy (anonymous):

whats the answer cuz I dont get it

OpenStudy (raden):

7203 + 1029 + 49 + 42 + 2 = ........ if still confuse, use the calculator

OpenStudy (anonymous):

8,325?

OpenStudy (anonymous):

is that the final answer

OpenStudy (raden):

yeah, correct!

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

can you help me with 1 more

OpenStudy (raden):

sure, if i can :P

OpenStudy (anonymous):

Use the Factor Theorem to determine if x - 5 is a factor of f(x) = x^4 - 8x^3- 45x^2 + 524x - 1120. Yes, x - 5 is a factor of f(x). No, x - 5 is not a factor of f(x).

OpenStudy (anonymous):

its no right?

OpenStudy (raden):

to knowing it is right or not, lets check it f(x) = x^4 - 8x^3- 45x^2 + 524x - 1120 if (x-5) is one of the factor, then f(5) must be equal zero x^4 - 8x^3- 45x^2 + 524x - 1120 (5)^4 - 8(5)^3- 45(5)^2 + 524(5) - 1120 = 625 - 8(125) - 45(25) + 524(5) - 1120 = 625 - 1000 - 1125 + 2620 - 1120 = ... (this for you)

OpenStudy (anonymous):

true

OpenStudy (anonymous):

its no

OpenStudy (raden):

if = 0, yes it is true if not equal 0, its not true

OpenStudy (anonymous):

it equals 0

OpenStudy (raden):

then say its "TRUE"

OpenStudy (anonymous):

is that right though

OpenStudy (anonymous):

that Yes, x - 5 is a factor of f(x).?

OpenStudy (raden):

yeah, i got calculation above equal 0 then Yes, x - 5 is a factor of f(x)

OpenStudy (anonymous):

ok cool thanks

OpenStudy (raden):

you're welcome

OpenStudy (anonymous):

:)